Please help me to answer this question "What is x and y for the term 9x+y=10 and x+4y=5 using the substitution?" And please teach how to solve it step by step... Please... <<>>

Answer by praseena(37)


Solution:
Here the equations are,
9x+y=10 and
x+4y=5
method of substitution:
consider the first equation,(also you can consider the second equation)
let us tAke the first one
9x+y=10
subtract 9x from both sides.
==> 9x + y - 9x = 10 - 9x
==> y = 10 - 9x
here we got the value of 'y' in terms of x.
right?
now substitute this value of y in the second equation.
that is x + 4 * ( 10 - 9x ) = 5
x + 4 * 10 - 4 * 9x = 5

x + 40 - 36X = 5

Subtract 40 from both sides of the equation
==> x + 40 - 36x - 40 = 5 - 40
==> x - 36x = -35
==> -35x = -35
now divide both sides of the equation by -35
==> x = 1
understood?
now substitute this value of 'x' , in the expression, y = 10 - 9x
==> y = 10 - 9 * 1
==> y = 10 - 9
=1
so the solution for the equations in
x=1
y=1